# Slice Emittance Preservation and Focus Control in a Passive Plasma Lens

### v1

@@ Introduction
Over the past two decades, PBAs have demonstrated major improvements in energy gain [6, 7, 8, 9], energy efficiency [10, 11, 12, 13], energy spread [14, 11, 12], and acceleration stability and reproducibility [15, 11, 16, 17]. However, while challenging applications such as PBA-driven free-electron lasers (FELs) have been demonstrated for longer wavelengths (IR – XUV) [18, 19, 20, 21], the performance of “conventional” FEL facilities has not yet been matched; X-ray FELs and linear colliders require a smaller normalized beam emittance ( mmmrad, see End Matter).
In a blown-out plasma wake, the focusing gradients are linear and emittance preservation is possible [22, 23, 24]. The natural beam size in this regime is given by the matched beta function [25], , where is the beam energy, is the unit charge, is the plasma electron density, and is the vacuum permittivity; this is typically on the sub-mm to cm scale for GeV-class PBA beams, compared to tens of meters in RF accelerators. In transport, the strong focusing required to accommodate these small beta functions can cause chromatic aberration in the beam, yielding a relative emittance growth for a single focusing optic of (see End Matter), where is the distance between the beam waist and the optic and is the root-mean-square (RMS) relative energy spread. A small beam size typically also corresponds to a large divergence, which drives emittance growth even in free-space propagation [26]. Therefore, it is an advantage for the optic to be placed closer to the waist, a distance limited by its linear focusing strength, , where is the length of the optic. Plasma lenses could realize this advantage.
@@
In this Letter, we experimentally demonstrate slice-emittance preservation at sub-mmmrad emittances and focus control at cm-scale waist betas with an electron-beam-driven PPL, thereby strengthening the technique’s potential for use with high-brightness PBAs. This was achieved by operating with a highly underdense plasma ( cm-3), which decreased the demands on the driver to reach a full blowout and limited the change in particle energy to the few-per-mille level.
@@
The smallest measured projected (slice) waist beta was 5.60.1 mm (1.60.1 mm), a demagnification factor of the incoming waist beta of 44 (166), suggesting a minimum slice RMS beam size below 0.5 m. A larger imaging magnification of around -10 (maximum for the actual object-plane range) enabled these smaller waists to be measured. Moving the lens upstream and downstream by 15 mm primarily affected the downstream waist location, leaving the emittances and waist betas nearly unaffected. This resulted from the lens being positioned near the large incoming beam waist; the incoming beam optics varied only little between different lens positions. The achieved waist betas covered a range relevant for matching into a GeV-class PBA, which, together with the flexible waist positioning, could offer a way of optimizing optics matching.
@@
While shorter optics are theoretically less chromatic [56], the lower-density, longer PPL regime investigated here offers advantages over the higher-density regime [35, 56]. The lower plasma density reduces the bunch-density requirements on the driver bunch, enabling a witness-emittance-preserving blown-out wake and allowing a large fraction of the driver bunch to focus. Longer lenses also induce less synchrotron radiation aberration for higher beam energies than shorter lenses, given identical waist parameters [57, 58]. If spatially overlapping with the adjacent PBA plasma-density ramps, these long plasma lenses are perhaps more akin to independently tunable ramp extensions than separate optics [59, 42], and may also allow reduction of the ramp length, thereby reducing emittance growth [60].
@@
In summary, we have for the first time experimentally demonstrated slice-emittance preservation in a PPL, the tunability of the waist parameters, and reduced aberration compared to focusing using quadrupole magnets. While emittance growth was observed for part of the data, the sources were identified and can be mitigated. A low-density PPL has been shown to have clear advantages as a focusing section for matching into a PWFA stage. These findings represent a critical step toward compact, beam-quality-preserving optics in novel accelerators.

### v2

@@ Introduction
Over the past decades, PBAs have demonstrated major improvements in energy gain [6, 7, 8, 9], energy efficiency [10, 11, 12, 13], energy spread [14, 11, 12], and acceleration stability and reproducibility [15, 11, 16, 17]. However, while challenging applications such as PBA-driven free-electron lasers (FELs) have been demonstrated for longer wavelengths (IR – XUV) [18, 19, 20, 21], the performance of RF-based FEL facilities has not yet been matched; X-ray FELs and linear colliders require a smaller normalized beam emittance ( mmmrad, see End Matter).
In a blown-out plasma wake, the focusing gradients are linear and emittance preservation is possible [22, 23, 24]. The natural beam size in this regime is given by the matched beta function [25], , where is the beam energy, is the unit charge, is the plasma electron density, and is the vacuum permittivity; this is typically on the sub-mm to cm scale for GeV-class PBA beams, compared to tens of meters in RF accelerators. The strong focusing required to accommodate these small beams can cause deleterious chromatic aberration, yielding a relative emittance growth of (see End Matter), where is the distance between the beam waist and the optic and is the root-mean-square (RMS) relative energy spread. A small beam size usually corresponds to a large divergence, which drives emittance growth even in free-space propagation [26]. Therefore, minimizing the waist-optic distance is advantageous, a distance limited by the linear focusing strength, , where is the length of the optic. High strength simultaneously enables more compact optical systems, which currently drive space requirements of PBA experiments. Plasma lenses could realize these advantages.
@@
In this Letter, we experimentally demonstrate slice-emittance preservation at sub-mmmrad emittances and focus control at cm-scale waist betas with an electron-beam-driven PPL, using beams with three orders of magnitude higher brightness than earlier work [29]. This was achieved by operating with a highly underdense plasma ( cm-3), which decreased the demands on the driver to reach a full blowout and limited the change in particle energy to the per-mille level.
@@
The smallest measured projected (slice) waist beta was 5.60.1 mm (1.60.1 mm), a demagnification factor of the incoming waist beta of 44 (166), suggesting a minimum slice RMS beam size below 0.5 m. A larger imaging magnification of around -10 (maximum for the actual object-plane range) enabled these smaller waists to be measured. Moving the lens upstream and downstream by 15 mm primarily affected the downstream waist location, leaving the emittances and waist betas nearly unaffected. This resulted from the lens being positioned near the large incoming beam waist; the incoming beam optics varied little between different lens positions. The achieved waist betas covered a range relevant for matching into a GeV-class PBA, which, together with the flexible waist positioning, could offer a way of optimizing optics matching.
@@
While shorter optics are theoretically less chromatic [56], the lower-density, longer PPL regime investigated here offers advantages over the higher-density regime [35, 56]. The lower plasma density reduces the bunch-density requirements on the driver bunch, allowing a large fraction of it to focus. Longer lenses also induce less synchrotron radiation aberration than shorter lenses, given identical waist parameters [57, 58]. If spatially overlapping with the adjacent PBA plasma-density ramps, long plasma lenses are perhaps more akin to independently tunable ramp extensions than separate optics, which could improve the performance of the system as a whole [59, 42, 60].
@@
In summary, we have for the first time experimentally demonstrated slice-emittance preservation in a PPL, the tunability of the waist parameters, and reduced aberration compared to focusing using quadrupole magnets. While emittance growth was observed for part of the data, the sources were identified and can be mitigated. A low-density PPL has been shown to have clear advantages as a focusing section for matching into a PWFA stage. Shown compatibility with beams three orders of magnitude brighter than previous work [29] marks these findings as a critical step toward compact, beam-quality-preserving optics in novel accelerators.